# Classifying fusion rules of anyons or SymTFTs: A general algebraic formula for domain wall problems and quantum phase transitions

### v2

@@ Abstract
Anyon has a significant role in many aspects of contemporary physics, such as quantum phase transitions and topological orders[1, 2, 3, 4], and quantum information[5, 6, 7]. Gapped or renormalization group (RG) domain wall[8, 9, 10, 11, 12, 13, 14, 15] is a fundamental theoretical framework producing classifications of anyons realized in conformal field theories (CFTs) or the corresponding topological quantum field theories(TQFTs)[16]. They are also expected to provide the condensation rule of anyons[17] or hierarchical structure in topologically ordered states, typically in fractional quantum Hall states[18, 19, 20, 21, 22, 23] (see also a recent work [24]). There exist several related but slightly different versions of definitions, but the most fundamental and common concepts are some canonical relationships between fusion rules of anyons by a morphism in abstract algebra or a functor in category theory. In particular, based on the recent development in the corresponding symmetry topological field theory (SymTFT)[25, 26, 27, 28], the most reasonable formulation for the domain walls is the relationship through ring homomorphisms between fusion rules[29, 14, 30, 15, 31, 32, 33]. This relation itself is also a natural structure to obtain the functor in $\mathbb{C}$-linear category established in [34], where $\mathbb{C}$ is the field of complex numbers (For readers interested in related historical aspects, we note [35]). Hence, we restrict our attention to this formalism with emphasis on linear algebra, providing a straighforward quantum mechanical description, whereas there exist other interesting formalisms respecting bulk and defect RG flows. For example, we note that the other formalism has a significant benefit in studying the defect or boundary critical phenomena[36, 37, 38, 39, 40, 41, 42] with a connection to the anomaly inflow mechnism[43], whereas their precise relation to the functor in category theory is less clear at this stage  [^1] We thank Gerard Watts and Ingo Runkel clarifying this point..
@@
Figure 1: Picture of the transformation of anyons through a domain wall. The UV objects and structures are represented in black, and those of the IR are represented in blue. The red colour represents the homomorphism or domain wall transforming anyons. Because of the surjective property of the homomorphism, the number of types of anyon decreases under the application of the homomorphism. One can see related figures in [32, 33], but we do not assume the coset or level-rank duality structures in the present paper.
@@
In many cases, the existence of ring homomorphism $\rho:\mathbf{A}\rightarrow\mathbf{A^{\prime}}$ is assumed and its consequence has been studied where $\mathbf{A}=\{\alpha\}_{\alpha}$ and $\mathbf{A^{\prime}}=\{\alpha^{\prime}\}_{\alpha^{\prime}}$ are the corresponding ultraviolet (UV) and infrared (IR) fusion ring and $\alpha$ and $\alpha^{\prime}$ are their anyons or symmetry operators. We also note that we do not assume the chirality or orientation of $\mathbf{A}$ and $\mathbf{A^{\prime}}$ (see [45, 26] for related discussions). The homomorphism provides the transformation law of anyon $\rho(\alpha)=\sum_{\alpha^{\prime}}A_{\alpha}^{\alpha^{\prime}}\alpha^{\prime}$ (Fig. 1). However, one can notice a puzzle in this research direction: How can one construct $\rho$ or determine the coefficients $A_{\alpha}^{\alpha^{\prime}}$? If the UV and IR theories are related by the coset construction[46, 47, 48] or level-rank duality[49, 50, 51], one can solve this problem straightforwardly by mapping the problem to a domain wall problem in TQFTs and projecting out the particle at the domain wall[33]. Related to this, in several research communities, it is known that some correspondence between the massless renormalization group (RG) flows[52] and the coset construction in CFTs should hold [53, 54, 55, 56, 57, 58]. However, in general, the construction of homomorphism is a difficult problem and requires extensive categorical or combinatorial calculations[59, 60, 61, 62, 63]. We also note related studies on conformal interfaces as the analytical analog [64, 9, 65, 66, 67, 68, 69, 70, 71, 72, 26, 73, 74, 28].
@@
In this paper, we propose a general formula, Eq.(10), for constructing homomorphisms between a wider class of fusion rings. Technically, our method requires some elementary calculations in abstract algebra (or linear algebra) [^2] This fundamental fact in abstract algebra has been notified to the author from Gerard Watts and Ingo Runkel.. Hence, we only note a classic textbook[76]. This simplicity is in sharp contrast with other frameworks. The most fundamental assumption in this manuscript is the Verlinde formula[77] (in commutative fusion rings). The resultant formula for the homomorphism is also similar to the Verlinde formula, and it provides a possible form of homomorphism between fusion rings. For readers interested in mathematical or categorical formulation, we note that the formula provides algebraic data of a functor between $\mathbb{C}$-linear categories.
@@
The rest of the manuscript is organized as follows. In Sec. II, we introduce the general formula, Eq.(10), constructing ring homomorphisms between fusion rings. This is the main result of this paper, and the formula applies to any categorical framework if the corresponding functor is well-defined. In Sec. III, we discuss the more involved classification by conformal spin of anyons or associated sectors after constructing the homomorphism. The implications for the anomaly cancellation mechanism and related categorical notions are also discussed. We study the correspondence between the ideal structure in massless RG and module in the massive RG and its analogy to Nambu-Goldstone type phenomenology in Sec. III.1. In Sec. IV, we discuss the application of our method to more general settings, combining our method with orbifolding, extension, or nontrivial ring isomorphism. The results are summarized in Eq.(17) and Eq.(18). Sec. V is the concluding remark, and we summarize the consequent phenomenology and comment on the future research directions.
@@ II Classification of anyons or SymTFTs
  • There exist $n$ and $n^{\prime}$ types of simple objects in the UV and IR theories, respectively.
@@
In the following discussion, we distinguish the UV and IR theories by specifying the prime symbol ’ to the IR theory. We denote the algebraic objects in SymTFTs in CFTs or anyons in TQFTs by the Greek symbols $\alpha,\beta,$.... For simplicity, we assume that the theories are diagonal $A$-type model, and the matrix representation of the modular $S$ transformation as $\chi_{\alpha}(-1/\tau)=\sum_{\beta}\chi_{\beta}S_{\alpha,\beta}$ where $\chi$ is the chiral character and $\tau$ is the modular parameter. In this setting, the modular invariant is $\sum_{\alpha}|\chi_{\alpha}|^{2}$.
@@
where $I$ is the identity operator corresponding to the vacuum and $N$ is the fusion matrix $\alpha\times\beta=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\gamma$. In this setting, one can represent the anyons by idempotents, $\{e_{(\alpha)}\}$[34],
@@
For the latter discussion, we note the benefits of introducing the idempotents. The idempotent representation provides algebraic data of a ring homomorphism straightforwardly. It should be noted that the anyon basis provides the natural phenomenology or intuitions, but their classifications by homomorphism are difficult to read. This may be a reason why the explicit coefficient of the gapped domain wall has not been calculated in the literature, and the resolution of this difficulty is the main result of the present paper.
By using the idempotents, one can directly obtain homomorphisms, by specifying the preserved idempotents $\{e_{(\alpha_{\text{pr}})}\}$, as follows:
@@
By assuming the linearlity, $\rho(\alpha+\beta)=\rho(\alpha)+\rho(\beta)$, and the compativility of the product, $\rho(e_{(\alpha)}\times\rho_{(\beta)})=\rho(e_{(\alpha)})\times\rho(e_{(\beta)})$, the above mapping provides the ring homomorphism. For latter use, we note the set of the label of preserved idempotent as $\{(\alpha_{\text{pr}})\}=\mathbb{S}_{\rho}$ and the set of the UV and IR pairs as $\mathbb{S}_{\rho,\times}=\{(\alpha_{\text{pr}},\alpha^{\prime}_{\text{pr}})\}$.
@@
This form itself is analogous to the Verlinde formula. One can see related discussions in [30, 15, 31], but our method only requires a modular $S$ matrix of the UV and IR theories and is more general and concise. Moreover, one can easily detect the structure of an ideal as idempotent $\{e_{(\alpha)}\}_{\alpha\notin\mathbb{S}_{\rho}}$. Hence, by identifying the ideal as module appearing in the massive RG as in [91, 92, 32, 33], one can obtain the module of massive RG which is dual to the massless flow as $\{|\alpha\rangle\rangle\}_{\alpha\notin\mathbb{S}_{\rho}}$, where $|\alpha\rangle\rangle$ is the (smeard) Ishibashi state[93], satisfying the relation $\langle\langle\alpha|e^{i\pi\tau H_{CFT}}|\beta\rangle\rangle=\chi_{\alpha}(\tau)\delta_{\alpha,\beta}$ where $H_{CFT}$ is the CFT Hamiltonian, $\tau$ is the modular parameter and $\delta_{\alpha,\beta}$ is the Kronekker delta. By applying the linear transformation, one can obtain the corresponding representation of (smeared) Cardy states[91] as
@@
where $\{\mathcal{Q}_{\alpha}\}$ is the set of topological symmetry operators (or conserved charges) satisfying the same fusion rule of $\{\alpha\}$. One can see related discussion in [92, 99, 100, 26, 32, 101, 102, 103, 33]. For readers interested in the historical aspect of the correspondence between BCFT and massive RG, we note earlier works in massive integrable models [104, 105] and recent works on the entanglement spectrum[106, 107, 108](See also the historical remarks in [109, 26, 102]).
@@ III Preserved sectors and unbroken subalgebra
Whereas we have constructed the general class of ring homomorphisms, their properties (or physical implications) will depend on more detailed data or the setting of the theory, such as central charge and conformal spin of the objects. More precisely, the anomaly-free conditions of the unbroken algebraic structure, $\mathbf{A}_{\text{ub}}$ with $\mathbf{A}_{\text{ub}}\cap\text{Ker}\rho=\{0\}$, play fundamental roles in studying lattice or quantum field theoretic realizations. In mathematical or categorical terms, when $\mathbf{A}_{\text{ub}}$ coincides with the connected etale algebra of $\mathbf{A}\boxtimes\overline{\mathbf{A^{\prime}}}$, the anomaly-free condition will be satisfied. At the ring theoretic level, $\mathbf{A}_{\text{ub}}\in\mathbf{A}$ and $\mathbf{A^{\prime}}_{\text{ub}}=\rho\left(\mathbf{A}_{\text{ub}}\right)\in\mathbf{A^{\prime}}$ are identical, but when introducing the braiding properties or constraint from the conformal spin, they can be different. The condition to obtain the connected etale algebra is nothing but the welldefinedness of the wavefunction constructed from $\mathbf{A}_{\text{ub}}\boxtimes\overline{\mathbf{A^{\prime}}}_{\text{ub}}$, and can be interpreted as the Lieb-Schultz-Mattis type anomaly free condition[112, 113, 114]: The conformal spin $s$ of the condensed objects $\alpha\boxtimes\overline{\rho(\alpha)}$ satisfy relation $s_{\alpha\boxtimes\overline{\rho(\alpha)}}\in\mathbb{Z}/2$ where $\alpha$ is a simple obejct in $\mathbf{A}_{\text{ub}}$, $\mathbb{Z}$ is the set of integers. Depending on the chirality of the theories, the expression of conformal spin $s$ by conformal dimensions will be changed, such as $s=\pm h_{\alpha}\pm h_{\rho(\alpha)}$, This condition itself appears commonly in the studies on TQFTs, for example in [115, 45, 116]. Hence, we note several works studying the corresponding structure in CFTs or SymTFTs[117, 54, 114, 118, 119, 120, 121, 122, 28, 123]. More historically, the corresponding objects for unbroken group symmetry has been studied in the name integer spin simple current or discrete torsion[124, 125, 126, 127, 128, 129].
@@
Similar to the anomaly-free condition of an unbroken subalgebra, there will exist an anomaly-free condition of preserved idempotents or sectors. Interestingly, if one assumes the anomaly-free condition for every preserved idempotent to the flow $M(4,5)\rightarrow M(3,4)$, the homomorphism is unique[33]. Moreover, the flow inevitably changes the chirality of anyons as in the Alice ring[130, 131].

### v3

@@ Abstract
Anyon[1, 2, 3, 4] has a significant role in many aspects of contemporary physics (see also historical remarks in [5]), such as quantum phase transitions and topological orders[6, 7, 8, 9], and quantum information[10, 11, 12]. Gapped or renormalization group (RG) domain wall[13, 14, 15, 16, 17, 18, 19, 20, 21] is a fundamental theoretical framework producing classifications of anyons realized in conformal field theories (CFTs) or the corresponding topological quantum field theories(TQFTs)[22]. They are also expected to provide the condensation rule of anyons[23, 24] or hierarchical structure in topologically ordered states, typically in fractional quantum Hall states[25, 26, 27, 28, 29, 30] (see also a recent work [31]). There exist several related but slightly different versions of definitions, but the most fundamental and common concepts are some canonical relationships between fusion rules of anyons by a morphism in abstract algebra or a (tensor) functor in category theory. In particular, based on the recent development in the corresponding symmetry topological field theory (SymTFT)[32, 33, 34, 35], the most reasonable formulation for the domain walls is the relationship through ring homomorphisms between fusion rules[36, 37, 20, 38, 21, 39, 40, 41]. This relation itself is also a natural structure to obtain the functor in $\mathbb{C}$-linear category established in [42], where $\mathbb{C}$ is the field of complex numbers (For readers interested in related historical aspects, we note [43]). There exists a reason to formulate the domain walls between $\mathbb{C}$-linear theories with a connection to the fundamental problem in realizing fractional quantum Hall systems in experimental settings[44, 45]. The results in [44, 45] indicate that the bulk-edge correspondence [6, 22, 46] predicted from the thermal Hall conductance[47] should be defined up to some connectivity of models[48, 49, 50]. Consequently, the experimental realizations indicate the changes of chiral central charges between underlying CFTs realized in the edge modes of topological orders, and this change of central charge is prohibited without the $\mathbb{C}$-linear formulations[51, 52]. Regardless of their experimental importance, the $\mathbb{C}$-linear formulations of the domain walls between CFTs or TQFTs are relatively less familiar, and still in development[21, 40, 41, 53]. Moreover, corresponding to the complex number coefficients before anyonic objects, the simplicity or locality of objects can be changed in $\mathbb{C}$ linear theories. This is consistent with the anomaly inflow mechanism[54], which permits the existence of exotic particles in the domain wall phenomena.
Hence, we restrict our attention to the $\mathbb{C}$ linear formalism with emphasis on linear algebra, providing a straighforward quantum mechanical description, whereas there exist other interesting formalisms respecting bulk and defect RG flows. For example, we note that the other formalism has a significant benefit in studying the defect or boundary critical phenomena[55, 56, 57, 58, 59, 60, 61] with a connection to the anomaly inflow mechnism[54]. However, their precise relation to the (tensor) functor in category theory is less clear at this stage  [^1] We thank Gerard Watts and Ingo Runkel clarifying this point..
@@
Figure 1: Picture of the transformation of anyons through a domain wall. The UV objects and structures are represented in black, and those of the IR are represented in blue. The red colour represents the homomorphism or domain wall transforming anyons. Because of the surjective property of the homomorphism, the number of types of anyon decreases under the application of the homomorphism. The fundamental point is that the fusion and linear sum, $\times$ and ++ (or $\sum$), are compatible with the application $\rho$. One can see related figures in [40, 41], but we do not assume the coset or level-rank duality structures in the present paper.
@@
In most of the references studying the gapped domain walls, either $\mathbb{C}$-linear or not, the properties of ring homomorphism $\rho:\mathbf{A}\rightarrow\mathbf{A^{\prime}}$ are assumed and their consequences have been studied where $\mathbf{A}=\{\alpha\}_{\alpha}$ and $\mathbf{A^{\prime}}=\{\alpha^{\prime}\}_{\alpha^{\prime}}$ are the corresponding ultraviolet (UV) and infrared (IR) fusion ring and $\alpha$ and $\alpha^{\prime}$ are their anyons or symmetry operators. We also note that we do not assume the chirality or orientation of $\mathbf{A}$ and $\mathbf{A^{\prime}}$ (see [63, 33] for related discussions). The homomorphism provides the transformation law of anyon $\rho(\alpha)=\sum_{\alpha^{\prime}}A_{\alpha}^{\alpha^{\prime}}\alpha^{\prime}$ (Fig. 1). However, one can notice a puzzle in this research direction: How can one construct $\rho$ or determine the coefficients $A_{\alpha}^{\alpha^{\prime}}$? We remind that the ring homomorphism is a linear mapping compatible with $\times$ operations, i.e. $\rho(\alpha\times\beta)=\rho(\alpha)\times\rho(\beta)$, and it has been clarified that the exact algebraic construction of $\rho$ requires complicated calculations even in familiar models[40, 41]. If the UV and IR theories are related by the coset construction[64, 65, 66] or level-rank duality[67, 68, 69], one can solve this problem straightforwardly by mapping the problem to a domain wall problem in TQFTs and projecting out the particle at the domain wall[41]. Related to this, in several research communities, it is known that some correspondence between the massless renormalization group (RG) flows[70, 71, 72] and the coset construction in CFTs should hold [73, 74, 75, 76, 77, 78]. One can see the related discussions for lattice-gauge models in [79]. However, in general, the construction of a homomorphism is a difficult problem and requires extensive categorical or combinatorial calculations[80, 81, 82, 83, 84]. We also note related studies on conformal interfaces as the analytical analog [85, 15, 86, 87, 88, 89, 90, 91, 92, 93, 33, 94, 95, 35]. Historically, this research direction has been explored in the Kondo or Kane-Fisher problem (or related wire-junction problem) in condensed matter[96, 97, 98, 99, 100, 101, 102].
@@
In this paper, we propose a general formula, Eq.(12), for constructing homomorphisms between a wider class of fusion rings. Technically, our method requires some elementary calculations in abstract algebra (or linear algebra) [^2] This fundamental fact in abstract algebra has been notified to the author by Gerard Watts and Ingo Runkel.. Hence, we only note a classic textbook[104]. This simplicity is in sharp contrast with other frameworks. The most fundamental assumption in this manuscript is the Verlinde formula[105] (in commutative fusion rings) and the underlying modular $S$ properties of the theories. The resultant formula for the homomorphism is also similar to the Verlinde formula, and it provides a possible form of homomorphism between fusion rings. For readers interested in mathematical or categorical formulation, we note that the formula provides algebraic data of a functor between $\mathbb{C}$-linear categories.
@@
The rest of the manuscript is organized as follows. In Sec. II, we introduce the general formula, Eq.(12), constructing ring homomorphisms between fusion rings. This is the main result of this paper, and the formula applies to any categorical framework if the corresponding functor is well-defined. In Sec. III, we discuss the more involved classification by conformal spin of anyons or associated sectors after constructing the homomorphism. The implications for the anomaly cancellation mechanism and related categorical notions are also discussed. We study the correspondence between the algebraic structure, ideal in mathematics, in massless RG and module in the massive RG and its analogy to Nambu-Goldstone type phenomenology in Sec. III.1. Sec. IV is the concluding remark, and we summarize the consequent phenomenology and comment on the future research directions. In End Matter, we discuss the application of our method to more general settings following research directions in [33, 35, 53]. The results are summarized in Eq.(32) and Eq.(33).
@@ II Classification of anyons or SymTFTs
  • There exist $n$ and $n$’ types of simple objects in the UV and IR theories, respectively.
@@
In the following discussion, we distinguish the UV and IR theories by specifying the prime symbol ’ to the IR theory. We denote the algebraic objects in SymTFTs in CFTs or anyons in TQFTs by the Greek symbols $\alpha,\beta,$…. For simplicity, we assume that the theories are diagonal $A$-type model with the following modular invariant,
$Z(\tau)=\sum_{\alpha}|\chi_{\alpha}|^{2}$ (1)
where $\chi$ is the chiral character and $\tau$ is the modular parameter. The matrix representation of the modular $S$ transformation is $\chi_{\alpha}(-1/\tau)=\sum_{\beta}\chi_{\beta}S_{\alpha,\beta}$ and this plays the fundamental role in studying the fusion rings via the Verlinde formula, as we discuss below.
@@
where $I$ is the identity operator corresponding to the vacuum and $N$ is the fusion matrix $\alpha\times\beta=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\gamma$. In this setting, one can represent the anyons by idempotents, $e{(\alpha)}$}[42],
@@
For the Verlinde lines, one can introduce the corresponding idempotents as projections[42];
$e_{(\alpha)}=P_{\alpha}=\sum_{M,\overline{M}}\left(|\alpha,M\rangle\langle\alpha,M|\right)\otimes\left(\overline{|\alpha,\overline{M}\rangle}\overline{\langle\alpha,\overline{M}|}\right)$ (6)
where $M$ ($\overline{M}$) is the label of (anti)chiral descendant states, producing the (anti)chiral character $\chi_{\alpha}$ ($\overline{\chi_{\alpha}})$ by taking their traces. More generally, one can introduce similar chiral or antichiral symmetry operators[60, 119, 120, 35], and we discuss related involved problems resulting from the chiralities or braiding properties in Sec. III. However, the algebraic aspects are robust against such subtleties.
For the latter discussion, we note the benefits of introducing the idempotents. The idempotent representation provides algebraic data of a ring homomorphism straightforwardly. It should be noted that the anyon basis provides the natural phenomenology or intuitions, but their classifications by homomorphism are difficult to read. This may be a reason why the explicit coefficient of the gapped domain wall has not been calculated in the literature, and the resolution of this difficulty is the main result of the present paper. We also note that the Verlinde formula and related categorical notions appeared with a close connection to the boundary or defect CFT[121, 122] (see also reviews[123, 124, 125] and a textbook[126]). We note earlier references on the related subfactor theory[127, 128] and related recent reviews[129, 43]. However, we stress again that the $\mathbb{C}$-linear properties, the representation of fusion algebra as conserved charges, has been clarified in [42].
By using the idempotents, one can directly obtain homomorphisms by specifying the preserved idempotents $\{e_{(\alpha_{\text{pr}})}\}$, as follows:
@@
Because of the linearlity, $\rho(\alpha+\beta)=\rho(\alpha)+\rho(\beta)$, and the compativility of the product, $\rho(e_{(\alpha)}\times\rho_{(\beta)})=\rho(e_{(\alpha)})\times\rho(e_{(\beta)})$, the above mapping provides the ring homomorphism, compativility of ++ and $\times$ operations under the application $\rho$. For latter use, we note the set of the label of preserved idempotent as $\{(\alpha_{\text{pr}})\}=\mathbb{S}_{\rho}$ and the set of the UV and IR pairs as $\mathbb{S}_{\rho,\times}=\{(\alpha_{\text{pr}},\alpha^{\prime}_{\text{pr}})\}$. We represents $\mathbb{S}_{\rho}$ as linear subspace, pseudo-ring, or set of the labels depending on the contexts, and we denote the broken (or unpreserved) sectors as $\mathbb{S}_{\rho}^{\text{c}}=\{e_{(\alpha)}\}_{\alpha\notin\mathbb{S}_{\rho}}$.
@@
This form itself is analogous to the Verlinde formula but it contains both contributions of the UV and IR structures. One can see related discussions in [38, 21, 39], but our method only requires a modular $S$ matrix of the UV and IR theories and is more general and concise. More or less, the existing methods usually require additional constraints such as the properties of domain walls, and these constraints restrict possible domain walls very strongly. For example, it is impossible to connects two theories with different central charges in the original arguments on the gapped domain wall[14, 16]. Hence this formulation is insufficient to explain the controvercies of thermal Hall conductance[44, 45], experimental realization of anomaly inflow mechanism[54, 48]. In [21], the domain wall has been extended to a $\mathbb{C}$-linear functor, but still additional constraints from the so-called domain wall $S$-matrix appear. However, except for [21], the irrational coefficients before objects are usually not permitted (see [36, 38, 39], for example), and this prohibits one from studying existing simple RG flows, such as the RG flow from the tricritical Ising model to the Ising model, in their formalisms[40]. Hence, we propose Eq. (12) which requires the weaker assumptions, to test the validity of the ring homomorphism and related ideas in future studies.
In our formalism, one can easily detect the structure of an ideal as idempotent $\mathbb{S}_{\rho}^{\text{c}}=\{e_{(\alpha)}\}_{\alpha\notin\mathbb{S}_{\rho}}$. An ideal $\mathbf{I}(\subset\mathbf{A})$ is a linear subspace of $\mathbf{A}$ such as $\mathbf{A}\times\mathbf{I}=\mathbf{I}$, and $\mathbb{S}_{\rho}^{\text{c}}$ satisfy this condition (we note a classic paper[130] and its English translation[131]). For more detailed explanations, let us assume the perturbative realization of the massless flow $\rho$ as $H_{\text{UV}}+\lambda H_{\text{pert},\rho}$ where $H_{\text{UV}}$ is the Hamiltonian of the UV theoery, $H_{\text{pert},\rho}$ is the perturbation triggering the flow $\rho$, and $\lambda$ is the coupling constant. When identifying the massless flow as $\lambda>0$, there will exist the corresponding dual massive RG with $\lambda<0$. More symbolically, the decomposition of the UV theory, $\mathbf{A}=(\mathbf{A}/\mathbb{S}_{\rho}^{\text{c}})\oplus\mathbb{S}_{\rho}^{\text{c}}$ provides the decomposition of the low-energy sector as the quotient ring $\mathbf{A}/\mathbb{S}_{\rho}^{\text{c}}$ and the high-energy sector as ideal $\mathbb{S}_{\rho}^{\text{c}}$ in the massless flow $\rho$. By applying the Noether’s ring isomorphism theorem, one can identify $\mathbf{A^{\prime}}$ as $\mathbf{A}/\mathbb{S}_{\rho}^{\text{c}}$, and the low energy sectors correctly reproduce the consequence of $\rho$. Under the dual operation $\lambda\Rightarrow-\lambda$, the roles of high and low energy sector are exchaged.
[…]
@@
where $\mathcal{Q}_{\alpha}=\sum_{\beta}\frac{S_{\alpha,\beta}}{S_{I,\beta}}P_{\beta}$ is the topological symmetry operator (or conserved charge) satisfying the same fusion rule of $\{\alpha\}$. One can see related discussion in [136, 141, 142, 33, 40, 143, 144, 145, 41]. For readers interested in the historical aspect of the correspondence between BCFT and massive RG, we note earlier works in massive integrable models [132, 133] (See also the historical remarks in [146, 33, 144]). In condensed matter theory, the resultant correspondence between BCFT (or chiral CFT) and the bulk states of the topological orders can be seen in the Li-Haldane conjecture[147] as has been clarified in [148]. More traditionally, this can be regarded as the cut and gluing properties of TQFTs and underlying CFTs with a close analogy to the tunneling problem[8].
@@ III Preserved sectors and unbroken subalgebra
Whereas we have constructed the general class of ring homomorphisms, their properties (or physical implications) will depend on more detailed data or the setting of the theory, such as central charge and conformal spin of the objects. More precisely, the anomaly-free conditions of the unbroken subalgebra, $\mathbf{A}_{\text{ub}}\subset\mathbf{A}$ with $\mathbf{A}_{\text{ub}}\cap\text{Ker}\rho=\{0\}$, play fundamental roles in studying lattice or quantum field theoretic realizations. The condition $\mathbf{A}_{\text{ub}}\cap\text{Ker}\rho=\{0\}$ implements the ring isomorphism $\mathbf{A^{\prime}}_{\text{ub}}=\rho\left(\mathbf{A}_{\text{ub}}\right)=\mathbf{A}_{\text{ub}}$, and this defines the preserved or unbroken symmetry structures in the homomorphism $\rho$. The folding trick[100] is a standard method to map the problem of the homomorphism $\rho:\mathbf{A}\rightarrow\mathbf{A^{\prime}}$ to the related BCFT of the tensor product of the coupled model $\mathbf{A}\otimes\overline{\mathbf{A^{\prime}}}$. When the $\mathbb{Z}_{2}$ group structure, $\{I,\psi\}$ at UV and $\{I^{\prime},\psi^{\prime}\}$ at IR, is preserved, i. e. $\mathbf{A}_{\text{ub}}=\mathbb{Z}_{2}$, the extension of the $\mathbf{A}\otimes\overline{\mathbf{A^{\prime}}}$ by the $\mathbb{Z}_{2}$ group ring $\{I\otimes\overline{I^{\prime}},\psi\otimes\overline{\psi^{\prime}}\}(\subset\mathbf{A}\otimes\overline{\mathbf{A^{\prime}}})$ naturally encodes the data of RG flows[150, 15]. More generally, for unbroken group ring structures, related discussion can hold [63, 33]. Through the CFT/TQFT correspondence, this is nothing but a generalization of Lieb-Shultz-Mattis type arguments[151, 152] for gapless phases[153, 74, 154, 155]. One can see related discussions in [34].
More generally, when $\mathbf{A}_{\text{ub}}$ coincides with the connected etale algebra of $\mathbf{A}\boxtimes\overline{\mathbf{A^{\prime}}}$ where $\boxtimes$ is the Deligne product, the anomaly-free condition will be satisfied[63, 156, 33]. Because the connected etale algebra has braiding data outside of ring theoretic data, the classification (or the naturality of theories) can be more involved. For example, even when $\mathbf{A}_{\text{ub}}({\subset}\mathbf{A})$ and $\mathbf{A^{\prime}}_{\text{ub}}(=\rho\left(\mathbf{A}_{\text{ub}}\right)\subset\mathbf{A^{\prime}})$ are identical at the ring theoretic level, but after introducing the braiding properties or constraint from the conformal spin, the classifications can be different. It should also be noted that the systematic classification of the connected etale algebra of the chiral-chiral or chiral-antichiral product of CFTs is still in development. We note recent works on the unbroken subalgebraic structure in massless RGs[157, 158, 159, 160, 161, 162, 163]. There exists much progress in the understanding of the unbroken subalgebraic structures, but the role of the connected etale algebra (or integer spin nonsimple currents) has rarely been discussed.
@@
The condition to obtain the connected etale algebra is nothing but the welldefinedness of the wavefunction constructed from $\mathbf{A}_{\text{ub}}\boxtimes\overline{\mathbf{A^{\prime}}}_{\text{ub}}$, and can be interpreted as the Lieb-Schultz-Mattis type anomaly free condition[151, 152, 164]: The conformal spin $s$ of the condensed objects $\alpha_{\text{ub}}\boxtimes\overline{\rho(\alpha_{\text{ub}})}$ satisfy relation $s_{\alpha_{\text{ub}}\boxtimes\overline{\rho(\alpha_{\text{ub}})}}\in\mathbb{Z}/2$ where $\alpha_{\text{ub}}$ is a simple object in $\mathbf{A}_{\text{ub}}$, $\mathbb{Z}$ is the set of integers. Depending on the chirality of the theories, the expression of conformal spin $s$ by conformal dimensions $h$ will be changed, such as $s_{\alpha_{\text{ub}}\boxtimes\overline{\rho(\alpha_{\text{ub}})}}=\pm h_{\alpha_{\text{ub}}}\pm h_{\rho(\alpha_{\text{ub}})}$. This (half) integer condition, $s_{\alpha_{\text{ub}}\boxtimes\overline{\rho(\alpha_{\text{ub}})}}\in\mathbb{Z}/2$, itself appears commonly in the studies on TQFTs, for example in [165, 63, 156]. We also note several works studying the corresponding structure in CFTs or SymTFTs[153, 74, 164, 166, 167, 168, 169, 157, 35, 170]. More historically, the corresponding objects for unbroken group symmetry has been studied in the name integer spin simple current or discrete torsion[171, 172, 173, 174, 175, 176]. Whereas the (half) integral condition is evident from the values of conformal spin in the above discussions, the studies on their categorical formulations have been still in development.
@@
Similar to the anomaly-free condition of an unbroken subalgebra, there will exist an anomaly-free condition of preserved idempotents or sectors. Interestingly, if one assumes the anomaly-free condition for every preserved idempotent to the flow from the tricritical Ising model, $M(4,5)$, to the Ising model, $M(3,4)$, the homomorphism is unique[41]. The fusion rule of the tricritical Ising model can be expressed as tensor product of the Ising and Fibonnacci fusion rules $M(4,5)=\text{Ising}\otimes\text{Fibonacci}=\{I^{[1]},\psi^{[1]},\sigma^{[1]}\}\otimes\{I^{[2]},\tau^{[2]}\}$. The conformal dimensions of the objects are $h_{I^{[1]}\otimes I^{[2]}}=0,h_{\psi^{[1]}\otimes I^{[2]}}=3/2,h_{\sigma^{[1]}\otimes I^{[2]}}=7/16,h_{I^{[1]}\otimes\tau^{[2]}}=3/5,h_{\psi^{[1]}\otimes\tau^{[2]}}=1/10,h_{\sigma^{[1]}\otimes\tau^{[2]}}=3/80$. On the other hand, the conformal dimensions of the Ising model, $M(3,4)=\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$, are $\{0,1/2,1/16\}$. Hence, the condition of the half-integral condition is satisified in the flow $M(4,5)\rightarrow\overline{M(3,4)}$, with $\mathbb{S}_{\rho,\times}=\{(I^{[1]}\otimes I^{[2]},\overline{I^{\prime}}),(\psi^{[1]}\otimes I^{[2]},\overline{\psi^{\prime}}),(\sigma^{[1]}\otimes I^{[2]},\overline{\sigma^{\prime}})\}$. This flow inevitably changes the chirality of anyons as in the Alice ring[177, 178], and it is surprising that even the simplest flow shows such an exotic structure.